OpenStudy (anonymous):

How do I find the area of a face given that the large rectangle is 400in by 255, nose (square) = 55 in on one side, eyes (circle) have diameters of 21in, mouth (isoceles triangle) height=27in base=94in. (The nose,eyes, and mouth aren't included in the area)

hero (hero):

Area of Big Rectangle minus area of eyes, nose, and mouth

hero (hero):

Get it?

OpenStudy (anonymous):

Sorry lol I just saw your reply so I would take 400(255)- \[\pi\]r^2- 1/2(27)(94)? How would I find the area of the square given that only one side of the square is 55inches?

hero (hero):

Area of a square = s^2 = 55^2

OpenStudy (anonymous):

Ahh I see! i'll work it out from here! Thank you so much!